Kevin is 6 years older than Keisha. Together their ages equal four times Keishas age. How old is each? Write in algabraic sentance and solve.
Answer by stanbon(75887)

Kevin is 6 years older than Keisha. Together their ages equal four times Keishas age. How old is each? Write in algabraic sentance and solve.
--------------------------------
v = k + 6
------------------
v+k = 4k
----------------
Substitute or v and solve for "k".
k+6 + k = 4k
2k = 6
k = 3 (Keisha's age now)
------------
Since v = k+6, v = 3+6 = 9 (Kevin's age now)
==============================================
Cheers,
Stan H.